Above what Fe2 concentration will Fe(OH)2 precipitate from a buffer solution that has a pH of 8.62? The Ksp of Fe(OH)2 is 4.87×10-17.
Answer #1

pt = 8.62 esp = 4.87 x 10-17 M pH & pot = 14 & pot = 14- pH - 14- 8.62 = 5:38 OH - - apoH - 5.38 10 = 10 = 4.168x100 Expressi